# Hybrid Hidden Markov Model for Modeling Equity Excess Growth Rate Dynamics: A Discrete-State Approach with Jump-Diffusion

### 10 Mar 2026 v1

@@ Abstract.
Generating synthetic financial time series that preserve the statistical properties of real market data is essential for stress testing, risk model validation, and scenario design, yet existing approaches, from parametric models to deep generative networks, struggle to simultaneously reproduce the heavy-tailed distributions, negligible linear autocorrelation, and persistent volatility clustering that characterize empirical equity data. We propose a hybrid hidden Markov framework that discretizes continuous excess growth rates into Laplace quantile-defined market states and augments the regime-switching process with a Poisson-driven jump-duration mechanism to enforce realistic tail-state dwell times. Parameters are estimated by direct transition counting, avoiding the Baum-Welch EM algorithm entirely. We evaluate synthetic data quality using three complementary metrics: Kolmogorov-Smirnov and Anderson-Darling goodness-of-fit pass rates for distributional fidelity, and ACF mean absolute error for temporal structure preservation. Applied to ten years of SPY data across 1,000 simulated paths, the framework achieves KS and AD pass rates exceeding 97% and 91% in-sample and 94% out-of-sample (full calendar year 2025), partially reproducing the ARCH effect that standard regime-switching models miss. No single model dominates all quality dimensions: GARCH(1,1) reproduces volatility clustering more accurately but fails distributional tests (5.5% KS pass rate), while the standard HMM without jumps achieves higher distributional fidelity but cannot generate persistent high-volatility regimes. The proposed hybrid framework offers the best joint quality profile across distributional, temporal, and tail-coverage metrics among the parametric models evaluated. A Single-Index Model extension propagates the SPY factor path to a 424-asset universe, enabling scalable correlated synthetic path generation while preserving cross-sectional correlation structure.
@@ 1. Introduction
Generating synthetic financial time series that faithfully preserve the statistical properties of real market data is a central challenge in quantitative finance and data quality research. High-fidelity synthetic data are essential for stress testing risk models against plausible but unobserved market scenarios, validating portfolio optimization algorithms, and augmenting limited training sets for machine learning pipelines (Assefa et al., 2020; Jordon et al., 2022). Yet the quality requirements for financial synthetic data are unusually demanding: empirical equity excess growth rates exhibit three well-documented statistical regularities, namely heavy-tailed (leptokurtic) distributions, negligible linear autocorrelation in raw returns, and persistent volatility clustering, that any generative framework must simultaneously reproduce to be considered statistically faithful (Mandelbrot, 1963; Cont, 2001; Stenger et al., 2024). These domain-specific quality criteria, often called stylized facts, serve as the benchmark against which the fidelity of synthetic financial data must be evaluated.
@@
Existing generative approaches address these quality requirements from different angles but each falls short on at least one dimension. GARCH-family models (Engle, 1982; Bollerslev, 1986) capture conditional variance dynamics but do not explicitly represent discrete market regimes or sudden jumps in volatility. Stochastic-volatility and jump-diffusion models (Heston, 1993; Merton, 1976) enrich tail behavior but lack a mechanism for volatility persistence after jumps. Deep generative models, including recurrent neural networks (Hochreiter and Schmidhuber, 1997; Fischer and Krauss, 2018) and generative adversarial networks (Takahashi et al., 2019; Kwon and Lee, 2024; Yoon et al., 2019), can learn complex distributions from data but struggle to reproduce temporal dependence structures, particularly volatility clustering, unless the network architecture is specifically designed to capture them (Stenger et al., 2024). Hidden Markov models (HMMs) and related regime-switching frameworks offer a probabilistic structure for non-stationary market behavior through latent state dynamics (Rabiner and Juang, 1986), yet standard HMM specifications fail to generate persistent high-volatility regimes, leading to overly rapid reversion following extreme events and poor temporal quality in the synthetic output (Rydén et al., 1998; Bulla and Bulla, 2006). Bridging this gap requires a generative framework that preserves both distributional fidelity and temporal structure while remaining computationally scalable.
@@
We address this challenge by developing a hybrid HMM framework that discretizes continuous excess growth rates into quantile-based regimes and augments the resulting Markov process with a two-parameter Poisson jump-duration mechanism that forces the model to dwell in high-volatility tail states for empirically realistic durations, connecting gradual price evolution with sudden large moves identified by jump-detection methods (Au Yeung et al., 2020). By partitioning excess growth rates into quantile-defined states via a Laplace cumulative distribution function, we enable direct frequentist counting of regime transitions, bypassing the iterative Baum-Welch algorithm (and its sensitivity to starting values) entirely, making the framework scalable to a 424-asset synthetic data pipeline. We evaluate synthetic data quality using five complementary metrics: Kolmogorov-Smirnov (KS) and Anderson-Darling (AD) pass rates for distributional fidelity, Wasserstein-1 and Hellinger distances for effect-size quantification, and autocorrelation function mean absolute error (ACF-MAE) for temporal structure preservation, all reported with standard errors. Using ten years of SPY data (2014–2024) for training and 249 trading days (full calendar year 2025) for out-of-sample testing, we demonstrate KS and AD pass rates of 97.6% and 91.3% in-sample and 94.4% and 95.1% out-of-sample across 1,000 simulated paths, with volatility clustering preserved in both windows. A Single-Index Model factor extension (Sharpe, 1963) propagates the univariate SPY engine to the full 424-asset universe in a single generative pass, preserving cross-sectional correlation structure. We find that the jump-duration extension substantially improves the reproduction of persistent high-volatility regimes relative to standard regime-switching models, though the improvement is partial: approximately 24% of simulated paths contain jump episodes, shifting the ensemble ACF toward the empirical profile without fully matching it. The framework does not dominate all quality dimensions; GARCH(1,1) achieves lower ACF error, and the standard HMM without jumps attains higher distributional pass rates. The contribution is a balanced quality profile that avoids the severe failures each alternative exhibits on its weakest dimension.
@@
Figure 1. Empirical stylized facts of SPY daily excess growth rates (2014–2024). Panel (a) shows the leptokurtic marginal distribution; the Laplace fit substantially outperforms the Gaussian. Panel (b) confirms heavy tails via a normal Q-Q plot. Panels (c) and (d) contrast the near-zero autocorrelation of raw returns (consistent with the efficient markets hypothesis) against the persistent autocorrelation of absolute returns, motivating the jump-duration extension.
@@ 2. Related Work
Research on reproducing the canonical stylized facts of financial markets divides into two broad paradigms: bottom-up agent-based models, which derive return dynamics from the simulated interaction of heterogeneous agents, and top-down data-driven models, which fit statistical structures directly to observed return series. In the bottom-up tradition, stylized facts emerge from collective behavior rather than being assumed. Lux and Marchesi (Lux and Marchesi, 1999) showed that a two-type fundamentalist-chartist population reproduces heavy tails and volatility clustering as emergent properties, Cont and Bouchaud (Cont and Bouchaud, 2000) demonstrated that a simple random-graph herding structure alone is sufficient to generate fat-tailed return distributions, and Chen, Tan, and Zheng (Chen et al., 2015) showed that multi-level herding simultaneously reproduces sector structure and volatility clustering. The minority game of Challet and Zhang (Challet and Zhang, 1997) introduced the paradigm of bounded-rational agents competing via strategic coordination, a framework subsequently shown to produce non-Gaussian fluctuations, and Farmer and Joshi (Farmer and Joshi, 2002) demonstrated that standard trading strategies simultaneously deployed by many agents endogenously generate volatility clustering and excess kurtosis without any exogenous noise assumption. LeBaron (LeBaron, 2006) surveys this class of models systematically; Arthur (Arthur, 2021) provides a recent synthesis arguing that heterogeneous agent interactions give rise to phenomena inaccessible to equilibrium theory, and Farmer (Farmer, 2025) demonstrates that quantitative ABMs can produce time-series predictions competitive with mainstream macroeconomic models. While mechanistically compelling, agent-based models can be challenging to calibrate directly to a specific asset’s return history and can be computationally demanding to deploy as practical synthetic data generators. Top-down models take the complementary approach: rather than explaining the origins of stylized facts, they specify statistical structures that match how those facts appear in data. GARCH, jump-diffusion processes, HMMs, and deep generative architectures all belong to this family; the subsections below survey each in turn, as they form the direct context for the present work.
@@ 2.1. Stylized facts and the limitations of parametric models
The quantitative study of financial market dynamics has been shaped by a tension between tractable theoretical frameworks and the empirical complexity of observed data. The Black-Scholes-Merton option pricing model (Black and Scholes, 1973) and the geometric Brownian motion (GBM) paradigm embedded within it assume that log-returns are independently and identically distributed Gaussian random variables, producing asset prices that move smoothly (no sudden jumps) with fixed-size random fluctuations and thin tails that assign negligible probability to extreme moves. These properties carry well-known practical advantages: closed-form option prices, simple portfolio variance calculations, and minimal data requirements. However, the Gaussian diffusion paradigm conflicts with a broad set of empirical regularities documented well before the model’s widespread adoption. Mandelbrot (Mandelbrot, 1963) demonstrated that speculative price changes exhibit heavy tails far beyond Gaussian predictions, a property long characterized in statistics as leptokurtosis. Fama (Fama, 1970) reviewed evidence that raw return series display negligible linear autocorrelation, consistent with the efficient markets hypothesis. Yet Schwert (Schwert, 1989) and others documented that the magnitude of returns exhibits substantial positive autocorrelation decaying slowly over time: large price changes tend to be followed by large changes regardless of sign, a pattern referred to as volatility clustering. Cont (Cont, 2001) provides a systematic review confirming these regularities across asset classes, geographic markets, and sampling frequencies. Collectively, leptokurtosis, the absence of return predictability, and volatility clustering define the benchmark against which any model of market dynamics must be evaluated.
@@
The most influential class of models designed to address volatility clustering within the Gaussian framework is the autoregressive conditional heteroscedasticity (ARCH) family introduced by Engle (Engle, 1982), generalized by Bollerslev (Bollerslev, 1986) to the GARCH(p,q) specification in which the conditional variance depends on both past squared residuals and past conditional variances. GARCH models are parsimonious and widely used, but they have well-documented limitations. Heavy tails must be injected by choosing a fat-tailed noise distribution (e.g., Student-t); the model itself does not produce them. There is no representation of discrete market regimes or sudden jumps. Volatility persistence is controlled by a single decay parameter, so the model cannot distinguish between a calm market that slowly heats up and a sudden crash that lingers (Diebold and Inoue, 2001). Stochastic volatility models (Stein and Stein, 1991; Heston, 1993) model volatility as a hidden quantity that evolves randomly over time, addressing some of these shortcomings, but require computationally intensive estimation and their parameters do not directly correspond to observable market states. An alternative approach to modeling fat tails and discontinuous price behavior is the explicit introduction of Poisson-driven jumps into the asset price process. Merton (Merton, 1976) added a compound Poisson jump component to GBM, producing a Gaussian-mixture marginal distribution that directly addresses leptokurtosis, and Kou (Kou, 2002) extended this framework with asymmetric jump sizes to capture the empirical pattern that crashes tend to be sharper than rallies. Jump-diffusion models are particularly relevant in equity markets, where macroeconomic announcements, earnings surprises, and geopolitical events generate discrete large-magnitude price movements (Au Yeung et al., 2020). However, in their standard form, jump-diffusion models do not provide a mechanism for volatility persistence: the post-jump volatility immediately returns to its baseline level, absent a separate stochastic volatility component.
@@ 2.2. Hidden Markov models and regime-switching
Hidden Markov models (HMMs) provide a probabilistic framework for non-stationary time series in which the observed data are generated by a process that switches among a discrete set of latent states according to Markov dynamics. The Baum-Welch expectation-maximization algorithm for maximum-likelihood estimation of emission and transition parameters was introduced by Baum, Petrie, Soules, and Weiss (Baum et al., 1970); Rabiner and Juang (Rabiner and Juang, 1986) provided the influential tutorial that brought HMMs to widespread attention in engineering and applied science. Hamilton (Hamilton, 1989) applied the regime-switching HMM to macroeconomic time series, demonstrating that U.S. GNP growth is well described by a two-state model with distinct expansion and recession regimes, establishing the template for latent-state modeling of financial time series. The identification of time-varying market regimes has since become a central application of HMMs in finance, with uses ranging from stock return modeling and volatility forecasting (Rossi and Gallo, 2006) to stock trading strategies (Nguyen, 2018).
Rydén, Teräsvirta, and Åsbrink (Rydén et al., 1998) provided an influential systematic evaluation of HMM specifications against the stylized facts of daily return series, showing that simple Gaussian-emission HMMs with a small number of states could reproduce leptokurtosis and the absence of return autocorrelation, but failed to generate the observed volatility clustering; the ACF of absolute returns under standard HMMs decayed far too rapidly relative to empirical data. This diagnostic limitation identified the core structural gap that the present study addresses. Bulla and Bulla (Bulla and Bulla, 2006) advanced this line of inquiry by considering hidden semi-Markov models, in which the time spent in each state can follow a flexible distribution rather than the simple memoryless (geometric) decay of standard Markov chains, and showed that this extension substantially improves the reproduction of volatility clustering by allowing the model to remain in high-volatility states for realistic durations. A persistent practical limitation of both standard and semi-Markov HMM approaches is the computational burden of maximum-likelihood estimation via the EM algorithm, particularly when the number of states is large or the asset universe is broad: the Baum-Welch procedure requires iterative forward-backward passes and may converge to local maxima depending on initialization. The present study addresses this limitation by replacing EM with direct frequentist counting of transitions between quantile-defined states, an approach that is computationally trivial and free of initialization sensitivity.
@@ 2.3. Synthetic data generation and deep generative models
The generation of synthetic financial time series has attracted increasing attention both as a practical tool for augmenting limited data, testing risk models, and stress testing portfolios, and as a benchmark for evaluating generative model quality. Assefa et al. (Assefa et al., 2020) survey the opportunities, challenges, and pitfalls of synthetic data generation in finance, noting that stylized-facts reproduction is the primary criterion for evaluating statistical quality and that naive generative approaches tend to fail on at least one of the three core stylized facts. The rise of generative adversarial networks (GANs) prompted several investigations into whether deep learning architectures could learn these properties from data. Takahashi, Chen, and Tanaka-Ishii (Takahashi et al., 2019) showed that standard GAN architectures produce synthetic return sequences that visually resemble financial time series but fail to reproduce ACF structure precisely; Kwon and Lee (Kwon and Lee, 2024) provided a careful evaluation against the stylized-facts checklist, finding that while GANs match marginal distributions reasonably well, capturing temporal dependence structures, particularly volatility clustering, remains challenging unless the architecture is specifically designed to learn them. Foundation models for Markov jump processes (Berghaus et al., 2024) offer a complementary perspective in which large-scale pretrained inference networks can be applied to new process instances without retraining, but their application to financial time series generation remains an open research direction.
Against this backdrop, the hybrid HMM framework of the present study occupies a distinct position: it is explicitly designed around the structural properties known to generate stylized facts, uses an interpretable discrete-state representation that can be audited and stress-tested, and achieves strong statistical fidelity with minimal computational overhead. The empirical estimation procedure avoids the training instability and data requirements of deep generative models, while the jump-duration mechanism provides a principled solution to the volatility-clustering limitation of standard HMMs. These properties make the framework particularly well-suited for risk management and stress-testing pipelines, where interpretability and computational reproducibility are as important as statistical fidelity.

### 02 Apr 2026 v2

@@ Abstract.
Generating synthetic financial time series that preserve the statistical properties of real market data is essential for stress testing, risk model validation, and scenario design. Existing approaches struggle to simultaneously reproduce heavy-tailed distributions, negligible linear autocorrelation, and persistent volatility clustering. We developed a hybrid hidden Markov framework that discretized excess growth rates into Laplace quantile-defined states and augmented regime switching with a Poisson jump-duration mechanism to enforce realistic tail-state dwell times. Parameters were estimated by direct transition counting, bypassing the Baum-Welch EM algorithm and scaling to a 424-asset pipeline. Applied to ten years of daily equity data, the framework achieved high distributional pass rates both in-sample and out-of-sample while partially reproducing the volatility clustering that standard regime-switching models miss. No single model was best at everything: GARCH(1,1) better reproduced volatility clustering but failed distributional tests, while the standard HMM without jumps passed more distributional tests but could not generate volatility clustering. The proposed framework delivered the most balanced performance overall. For multi-asset generation, copula-based dependence models that preserved each asset’s marginal HMM distribution substantially outperformed a Single-Index Model factor baseline on both per-asset distributional accuracy and correlation reproduction.
@@ 1. Introduction
Generating synthetic financial time series that faithfully preserve the statistical properties of real market data is a central challenge in quantitative finance and data quality research. High-fidelity synthetic data are essential for stress testing risk models against plausible but unobserved market scenarios, validating portfolio optimization algorithms, and augmenting limited training sets for machine learning pipelines (Assefa et al., 2020; Jordon et al., 2022). Yet the quality requirements for financial synthetic data are demanding: empirical equity excess growth rates exhibit three well-documented statistical regularities, namely heavy-tailed (leptokurtic) distributions, negligible linear autocorrelation in raw returns, and persistent volatility clustering, in which large changes tend to be followed by large changes of either sign and small changes by small changes (Mandelbrot, 1963), that any generative framework must simultaneously reproduce to be considered statistically faithful. These domain-specific quality criteria, often called stylized facts, serve as the benchmark against which the fidelity of any synthetic financial data must be evaluated (Figure 1) (Mandelbrot, 1963; Cont, 2001; Stenger et al., 2024). All three regularities are visible in SPY daily excess growth rates (Figure 1): the heavy-tailed distribution (1a), departures from normality in the tails (1b), near-zero autocorrelation of raw returns (1c), and the slow decay of absolute-return autocorrelation that characterizes volatility clustering (1d). The challenge of reproducing all three stylized facts simultaneously has been a long-standing problem in financial time series generation, and no single generative framework has yet succeeded in doing so at scale.
@@
Existing generative approaches each excel at one stylized fact but fail at others. GARCH-family models (Engle, 1982; Bollerslev, 1986) capture conditional variance dynamics but do not represent discrete regimes; stochastic-volatility and jump-diffusion models (Heston, 1993; Merton, 1976) enrich tail behavior but lack volatility persistence after jumps; deep generative models (Hochreiter and Schmidhuber, 1997; Fischer and Krauss, 2018; Takahashi et al., 2019; Kwon and Lee, 2024; Yoon et al., 2019) can learn complex distributions but struggle to reproduce volatility clustering (Stenger et al., 2024). Hidden Markov models (HMMs) offer a probabilistic regime-switching structure (Rabiner and Juang, 1986), yet standard HMMs fail to generate persistent high-volatility regimes, reverting too quickly after extreme events (Rydén et al., 1998; Bulla and Bulla, 2006). Bridging this gap requires a framework that preserves both distributional fidelity and temporal structure while remaining computationally scalable.
@@
In this study, we addressed this challenge by developing a hybrid HMM framework that discretized continuous excess growth rates into quantile-based regimes. The resulting Markov process was augmented with a two-parameter Poisson jump-duration mechanism that forced the model to dwell in high-volatility tail states for empirically realistic durations. By partitioning excess growth rates into quantile-defined states via a Laplace cumulative distribution function, we enabled direct frequentist counting of regime transitions, bypassing the iterative Baum-Welch algorithm (and its sensitivity to starting values) entirely, making the framework scalable to a 424-asset synthetic data pipeline. We evaluated synthetic data quality using five complementary metrics: Kolmogorov-Smirnov (KS) and Anderson-Darling (AD) pass rates for distributional fidelity, Wasserstein-1 and Hellinger distances to measure how far synthetic distributions deviated from the observed data, and autocorrelation function mean absolute error (ACF-MAE) for temporal structure preservation. Using a dataset of 424 United States-listed equities spanning ten years (2014–2024) for training and one year for out-of-sample testing, with the SPDR S&P 500 ETF Trust (SPY) as the primary validation asset, the framework achieved high distributional pass rates both in-sample and out-of-sample while partially reproducing the volatility clustering that standard regime-switching models miss. A Single-Index Model factor extension (Sharpe, 1963) propagated the univariate SPY engine to the full 424-asset universe, preserving cross-sectional correlation structure but producing poor distributional fits for assets that did not closely track the index. To address this limitation, we evaluated copula-based dependence models (Gaussian, Student-t, and C-vine) that preserved each asset’s fitted HMM marginal while injecting cross-asset correlation via rank reordering (Embrechts et al., 2002; Aas et al., 2009); the Student-t copula substantially outperformed the linear factor baseline on both per-asset distributional accuracy and correlation reproduction. Tunable hyperparameters controlled the jump frequency and duration, allowing users to trade off temporal and distributional quality for a given asset. No single model was best at everything: GARCH(1,1) better reproduced volatility clustering but failed distributional tests, and the standard HMM without jumps passed more distributional tests but could not generate volatility clustering. Thus, the proposed framework avoided the failures of each alternative and delivered the most balanced performance overall.
@@
Figure 1. Empirical stylized facts of SPY daily excess growth rates (2014–2024). Panel (a) shows the leptokurtic marginal distribution; the Laplace fit substantially outperforms the Gaussian. Panel (b) confirms heavy tails via a normal quantile-quantile (Q-Q) plot. Panels (c) and (d) contrast the near-zero autocorrelation of raw returns (consistent with the efficient markets hypothesis) against the persistent autocorrelation of absolute returns, motivating the jump-duration extension.
@@ 2. Related Work
Research on reproducing the canonical stylized facts of financial markets divides into bottom-up agent-based models, which derive return dynamics from simulated interactions of heterogeneous agents (Lux and Marchesi, 1999; Cont and Bouchaud, 2000; Chen et al., 2015; Farmer and Joshi, 2002; LeBaron, 2006; Arthur, 2021; Farmer, 2025), and top-down data-driven models, which fit statistical structures directly to observed return series. While agent-based models produce stylized facts as emergent properties, they can be challenging to calibrate to specific assets and computationally demanding as practical generators. The data-driven tradition is the focus of the present review; the subsections below trace the evolution from parametric return models through latent-state and deep generative approaches to the multi-asset extensions that motivate this study.
@@ 2.1. Stylized facts and the limitations of parametric models
The Black-Scholes-Merton option pricing model (Black and Scholes, 1973) and its underlying geometric Brownian motion (GBM) assume that log-returns are i.i.d. Gaussian: prices move smoothly, volatility is constant, and extreme moves are negligibly rare. These properties carry well-known practical advantages: closed-form option prices, simple portfolio variance calculations, and minimal data requirements. However, the Gaussian diffusion paradigm conflicts with a broad set of empirical regularities documented well before the model’s widespread adoption. Mandelbrot (Mandelbrot, 1963) demonstrated that speculative price changes exhibit heavy tails (leptokurtosis) beyond Gaussian predictions. Fama (Fama, 1970) reviewed evidence that raw return series display negligible linear autocorrelation, consistent with the efficient markets hypothesis. Yet Mandelbrot also observed that large price changes tend to cluster together regardless of sign, a pattern now called volatility clustering. Schwert (Schwert, 1989) and others later confirmed this statistically, showing that the autocorrelation of absolute returns decays slowly over time. Cont (Cont, 2001) provides a systematic review confirming these regularities across asset classes, geographic markets, and sampling frequencies. Collectively, heavy tails (leptokurtosis), the absence of return predictability, and volatility clustering define the benchmark against which any model of market dynamics must be evaluated.
@@
Engle (Engle, 1982) introduced autoregressive conditional heteroscedasticity (ARCH) models to address volatility clustering, and Bollerslev (Bollerslev, 1986) generalized this to the GARCH(p,q) specification. GARCH models are widely used but have well-documented limitations. Heavy tails must be injected by choosing a fat-tailed noise distribution (e.g., Student-t); the model itself does not produce them. There is no representation of discrete market regimes or sudden jumps. Volatility persistence is controlled by a single decay parameter, so the model cannot distinguish between a calm market that slowly heats up and a sudden crash that lingers (Diebold and Inoue, 2001). Stochastic volatility models (Stein and Stein, 1991; Heston, 1993) treat volatility as a latent process, but require computationally intensive estimation. An alternative approach introduces Poisson-driven jumps directly into the price process. Merton (Merton, 1976) added a compound Poisson jump component to GBM, producing a Gaussian-mixture marginal distribution that directly addresses leptokurtosis, and Kou (Kou, 2002) extended this framework with asymmetric jump sizes to capture the empirical pattern that crashes tend to be sharper than rallies. Jump-diffusion models are particularly relevant in equity markets, where macroeconomic announcements, earnings surprises, and geopolitical events generate discrete large-magnitude price movements (Au Yeung et al., 2020). However, in their standard form, jump-diffusion models do not provide a mechanism for volatility persistence: the post-jump volatility immediately returns to its baseline level, absent a separate stochastic volatility component. Hidden Markov models offer a structurally different solution by embedding these dynamics in a latent discrete-state process.
@@ 2.2. Hidden Markov models and regime-switching
Hamilton (Hamilton, 1989) introduced the regime-switching HMM to economics, showing that U.S. GNP growth is well described by a two-state model with distinct expansion and recession regimes. Since then, HMMs have become a standard tool for latent-state modeling of financial time series, with applications ranging from volatility forecasting (Rossi and Gallo, 2006) to equity trading strategies (Nguyen, 2018). Rydén, Teräsvirta, and Åsbrink (Rydén et al., 1998) systematically evaluated HMMs against the stylized facts of daily returns. Gaussian-emission HMMs with few states reproduce heavy tails and the absence of return autocorrelation, but fail to capture volatility clustering: the ACF of absolute returns decays far too quickly relative to empirical data. Bulla and Bulla (Bulla and Bulla, 2006) showed that hidden semi-Markov models (HSMMs), which replace the geometric sojourn-time distribution with a flexible alternative, substantially improve volatility-clustering reproduction by allowing realistic persistence in high-volatility states. However, both standard and semi-Markov approaches rely on EM estimation (Baum-Welch (Baum et al., 1970)), which requires iterative forward-backward passes and is sensitive to initialization. The present study sidesteps this limitation by replacing EM with direct frequentist counting of transitions between quantile-defined states, an approach that is computationally trivial and initialization-free. Alongside HMM-based methods, a parallel line of work has explored whether deep generative models can learn stylized facts directly from data.
@@ 2.3. Synthetic data generation and deep generative models
Synthetic financial time series are increasingly used to augment limited data, stress test portfolios, and benchmark generative models. Assefa et al. (Assefa et al., 2020) surveyed this landscape, noting that stylized-facts reproduction is the primary quality criterion and that naive generators tend to fail on at least one core stylized fact. Deep generative models, particularly generative adversarial networks (GANs), have attracted interest as a way to learn these properties directly from data, but results have been mixed. Takahashi, Chen, and Tanaka-Ishii (Takahashi et al., 2019) found that standard GAN architectures produced return sequences that visually resembled financial data yet failed to reproduce ACF structure, and Kwon and Lee (Kwon and Lee, 2024) confirmed that volatility clustering remained elusive unless the architecture was specifically designed for it. Foundation models for Markov jump processes (Berghaus et al., 2024) offer a complementary pretrained-inference perspective, though their application to financial series remains open. The hybrid HMM of the present study takes a different path: it replaces iterative EM with direct frequentist counting (eliminating training instability) and augments the Markov chain with a jump-duration mechanism that addresses the volatility-clustering gap without the architectural complexity of deep generative models. All of the models discussed so far are univariate; extending them to multi-asset portfolios introduces a separate set of challenges.